OpenStudy (anonymous):

Find the general solution for cos x = 0, where n is an integer.

OpenStudy (mertsj):

Remember that the cosine is like x. So the cosine is 0 when the angle is 90 degrees or 270 degrees. Every trip around the circle will have cosine of 0 at those points so the solution is : \[x=90, 270 +2\pi n\]

OpenStudy (mertsj):

If you want it in terms of radians, it is: \[x= \frac{\pi}{2}, \frac{3\pi}{2} + 2\pi n\]

OpenStudy (mertsj):

The one in degrees should have been +360 n

OpenStudy (anonymous):

can u help me with another one 3csc x=6

OpenStudy (anonymous):

would the answer be pi/6 + 2pi n

OpenStudy (anonymous):

and 5pi/6 + 2pi n

OpenStudy (mertsj):

\[\csc x=2\] \[\sin x=\frac{1}{2}\] \[x=\frac{\pi}{6}, \frac{5\pi}{6}+2\pi n\]

OpenStudy (mertsj):

Good for you. You got that one.

OpenStudy (anonymous):

and for this one ten= \[\sqrt{3}\]

OpenStudy (anonymous):

tan=\[\sqrt{3}\]

OpenStudy (anonymous):

is the answer pi/3 + npi

OpenStudy (mertsj):

\[x=\frac{\pi}{3}, \frac{4\pi}{3} + 2\pi n\]

OpenStudy (mertsj):

Yes. You could just say pi/3 +n pi. It would be the same thing.

OpenStudy (anonymous):

and 2 more question: i just want to make sure

OpenStudy (mertsj):

ok

OpenStudy (anonymous):

If you wanted to solve the equation 2sin x = 1, then you could find the solutions by looking at

OpenStudy (anonymous):

here are the options: the x-intercepts of y = 2sin x - 1 the y-intercepts of 2sin x = 1 the y-intercepts of y = 2sin x - 1 the x-intercepts of 2sin x = 1

OpenStudy (mertsj):

The first one.